Escuha Company produces two type of calculators: scientific and business. Both products pass through two producing departmentRequired: 1. Compute the overhead cost per unit for each product using a plantwide, unit-based rate using direct labor hours.
Answer #1
Escuha Company
Answer 1
Calculation of Plant wide OH Rate Department 1 Department 2 Total Note
Overhead costs    894,500.00       791,000.00 1,685,500.00 A
Direct Labor Hours    186,630.00       927,670.00 1,114,300.00 B
Predetermined OH Rate                   1.51 C=A/B
Overhead cost per unit Scientific Business Note
Direct Labor Hours    101,300.00    1,013,000.00 D
Predetermined OH Rate                1.51                   1.51 See C
Allocated OH 152,963.00 1,529,630.00 E=C*D
Units produced      72,000.00       720,000.00 F
Overhead cost per unit                2.12                   2.12 G=E/F
Answer 2
Calculation of Departmental OH Rate Department 1 Department 2
Overhead costs    894,500.00       791,000.00 See A
Direct Labor/ Machine Hours    382,120.00       927,670.00 H
Predetermined OH Rate                2.34                   0.85 I=A/H
Overhead cost per unit Scientific Business
Machine Hours      23,300.00       358,820.00 J
Predetermined OH Rate- Dept. 1                2.34                   2.34 See I
Allocated OH- Dept. 1      54,522.00       839,638.80 K=J*I
Labor Hours      26,100.00       901,570.00 L
Predetermined OH Rate- Dept. 1                0.85                   0.85 See I
Allocated OH- Dept. 2      22,185.00       766,334.50 M=L*I
Total Allocated OH      76,707.00 1,605,973.30 N=K+M
Units produced      72,000.00       720,000.00 See F
Overhead cost per unit                1.07                   2.23 O=N/F
Calculation of Departmental OH Rate Department 1 Department 2
Overhead costs    894,500.00       791,000.00 See A
Direct Labor/ Machine Hours    186,630.00       139,280.00 P
Predetermined OH Rate                4.79                   5.68 Q=A/P
Overhead cost per unit Scientific Business
Labor Hours      75,200.00       111,430.00 R
Predetermined OH Rate- Dept. 1                4.79                   4.79 See Q
Allocated OH- Dept. 1 360,208.00       533,749.70 S=R*Q
Machine Hours      24,100.00       115,180.00 T
Predetermined OH Rate- Dept. 1                5.68                   5.68 See Q
Allocated OH- Dept. 2 136,888.00       654,222.40 U=Q*T
Total Allocated OH 497,096.00 1,187,972.10 V=S+U
Units produced      72,000.00       720,000.00 See F
Overhead cost per unit                6.90                   1.65 W=V/F
Answer 3
Calculation of ABC rates for overhead X Y Z=X+Y AA AB AC=AB+AC AD=Z/AC
Activity cost pool Cost Driver Department 1 Department 2 Cost Scientific Business Quantity ABC rates
Setups Production runs       227,500.00       227,500.00       455,000.00            100.00               150.00          250.00 1,820.00
Inspection Inspection hours       190,000.00       190,000.00       380,000.00         2,000.00            3,000.00       5,000.00        76.00
Power Machine Hours       268,000.00       155,000.00       423,000.00      47,400.00       474,000.00 521,400.00          0.81
Maintenance Maintenance hours       209,000.00       218,500.00       427,500.00         2,250.00            9,000.00     11,250.00        38.00
Total Cost assigned 1,685,500.00
See AD See AA AE=AD*AE See AB AF=AD*AB
Allocation table Scientific
Business
Activity cost pool Cost Driver ABC Rates Activity used Cost assigned Activity used Cost assigned
Setups Production runs            1,820.00               100.00       182,000.00            150.00       273,000.00
Inspection Inspection hours                 76.00            2,000.00       152,000.00         3,000.00       228,000.00
Power Machine Hours                   0.81         47,400.00         38,454.55    474,000.00       384,545.45
Maintenance Maintenance hours                 38.00            2,250.00         85,500.00         9,000.00       342,000.00
Total Cost assigned       457,954.55 1,227,545.45 AG
Number of units         72,000.00       720,000.00 See F
Overhead cost per unit                   6.36                   1.70 AH=AG/F